•The difference between two numbers is 27. When twice the smaller number is subtracted from three times the larger number the difference is 95. What are the numbers?
Answer by Maths68(1474)

Let
Larger Number = x
Smaller Number = y
The difference between two numbers is 27
x-y=27
x=27+y............(1)
When twice the smaller number is subtracted from three times the larger number the difference is 95
2y-3x=95............(2)
Put the value of x in above equation
2y-3(27+y)=95
2y-81-3y=95
-y=95+81
-y=176
y=-176
Put the value of y in (1)
x=27-y............(1)
x=27-176
x=-149
Larger Number = -149
Smaller Number = -176



Check
Put the values of x and y in (1)
x=27+y............(1)
-149=27-176
-149=-149
Now Put the values of x and y in (2)
2y-3x=95............(2)
2(-176)-3(-149)=95
-352+447=95
95=95